Fórmulas Tasa de Fallas CGE Consultores

René González Leiva

2173037-8

Estudiante Ingeniería Civil Informática. Septiembre, 2008

1. Formulas

λ(t) = \frac{β}{α}×(\frac{t}{α})^{β−1}  (1)
λ(t) = cte  (2)

λ(t) = \frac{R(t − Δt) − R(t)}{R(t − Δt) × Δt}  (3)